OpenStudy (helpmeplease14):

WILL GIVE MEDAL IF HELP!!!! so there's a triangle. The width says 5. The other two sides are x-4 and x-5. The next triangle's width is 16. The other two sides are 10 and x-10. It's asking me to find the value of x so that the triangles can have the same perimeter. Idk what to do! HELP!!!!

OpenStudy (campbell_st):

well to have the same perimeters triangle 1 \[P_{tri~1} = 5 + x - 4 + x - 5\] you need to simplify this triangle 2 \[P_{tri ~2} = 16 + 10 + x - 10\] simplify this then its \[P_{tri~1} = P_{tri~2}\] so let the 2 perimeter equations equal each other and solve for x

OpenStudy (helpmeplease14):

Idk how to do it😢

OpenStudy (campbell_st):

ok.. can you simplify the 1st triangle perimeter P = 5 + x - 4 + x - 5

OpenStudy (helpmeplease14):

Okay so I know x are like term so...

OpenStudy (helpmeplease14):

It'll be 2x

OpenStudy (helpmeplease14):

Right?

OpenStudy (campbell_st):

almost P = 2x - 4 now simplify triangle 2

OpenStudy (helpmeplease14):

Ok

OpenStudy (helpmeplease14):

16+x

OpenStudy (campbell_st):

great so now it says the perimeters need to be the same... which means equal so 2x - 4 = x + 16 now just solve for x hope it helps

OpenStudy (helpmeplease14):

Then it'll be 2x-4=16-x

OpenStudy (helpmeplease14):

Tysm